# UFGraphFR: Graph Federation Recommendation System based on User Text description features

### v3

@@ 5.2.2 Implementation details
We implemented this method based on the Pytorch framework. In addition, we developed a variant of the method based on the original model, named UFGraphFR-Lite. UFGraphFR-Lite can improve operational efficiency by periodically updating user diagrams.
@@
Table 2: Performance comparison on four datasets.
@@
    Performance comparison of the four datasets. The best results are in bold and the best baseline results are underlined. FedPerGNN could not run on HetRec2011 due to unacceptable memory allocation (indicated by ‘-’). “ to indicate this). “*” and ‘Improvement’ denote statistically significant improvement and performance improvement over the best baseline, respectively. $\uparrow$ denotes the difference in enhancement relative to the best baseline
@@ 5.4 Ablation experiment (Q2)
Table 3: Ablation study results of UFGraphFR on four datasets. GPFedRec is used as the baseline. Performance drops relative to the full model are shown below each variant.
@@
GPFedRec (baseline) 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
UFGraphFR (w/o Transformer) 72.85 44.73 72.80 43.84 83.94 75.66 69.33 43.37
@@
UFGraphFR (Full) 76.03 47.31 75.55 46.32 85.69 77.32 70.42 45.17
    $\downarrow$ denotes the difference in performance degradation relative to the full model of UFGraphFR. $\uparrow$ denotes the difference in enhancement relative to the baseline model GPFedRec. Bolding is the highest performance for the current dataset
As shown in Figure 2, in our experiments, the model that removes the Sequence Modeling module (w/o Transformer) removes the Transformer block layer on the right side; the model that removes the Joint Embedding module (w/o Joint Embedding) removes the left side of the Joint Embedding is replaced with a normal Embedding layer. Therefore, the model that removes both modules (w/o Both) is to implement both of the above implementations simultaneously. However, compared with the baseline model GPFedRecLABEL:zhang_gpfedrec_2024, it still has an additional UserMLP layer for further learning user preferences.
5.4.1 w/o Transformer
@@ 5.5 PLM & Modeling Effect(Q3)
  • 2.
    LaBSE[51] The language-agnostic BERT sentence embedding encodes text into high-dimensional vectors. The Embedding dimension is 768.
  • 3.
    GTR-T5-Base,T5[52] This is a sentence-transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space.
  • 4.
    all-MiniLM-L6-v2,MiniLM This is a sentence-transformers model: It maps sentences & paragraphs to a 384 dimensional dense vector space and can be used for tasks like clustering or semantic search.
@@ 6 Conclusion
7 Declaration of Generative AI
During the preparation of this work the author(s) used Deepseek and ChatGPT in order to polished language. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v4

@@ 5.2.2 Implementation details
We implemented this method on the basis of the Pytorch framework. In addition, we developed a variant of the original model-based method, named UFGraphFR-Lite. UFGraphFR-Lite can improve operational efficiency by periodically updating user diagrams.
@@
Table 4: Performance comparison on four datasets. 00footnotetext: Performance comparison of the four datasets. The best results are in bold and the best baseline results are underlined. FedPerGNN could not run on HetRec2011 due to unacceptable memory allocation (indicated by ‘-’). “ to indicate this). “*” and ‘Improvement’ denote statistically significant improvement and performance improvement over the best baseline, respectively. $\uparrow$ denotes the difference in enhancement relative to the best baseline
@@ 5.4 Ablation experiment (Q2)
Table 5: Ablation study results of UFGraphFR on four datasets. GPFedRec is used as the baseline. Performance drops relative to the full model are shown below each variant. 00footnotetext: $\downarrow$ denotes the difference in performance degradation relative to the full model of UFGraphFR. $\uparrow$ denotes the difference in enhancement relative to the baseline model GPFedRec. Bolding is the highest performance for the current dataset.Among them, the variant UFGraphFR-nT(no Transformer) represents the version without Transformer; UFGraphFR-nJ(no Jointed-Embedding) represents the version without Jointed Embedding; UFGraphFR-nB(no Both) represents the version with both removed;
@@
GPFedRec 72.85 43.77 72.17 43.61 83.44 74.11 69.41 43.34
UFGraphFR-nT
72.85 44.73 72.80 43.84 83.94 75.66 69.33 43.37
@@
UFGraphFR 76.03 47.31 75.55 46.32 85.69 77.32 70.42 45.17
As shown in Figure 3, in our experiments, the model that removes the Sequence Modeling module (UFGraphFR-nT) removes the Transformer block layer on the right side; the model that removes the Joint Embedding module (UFGraphFR -nJ) removes the left side of the Joint Embedding is replaced with a normal Embedding layer. Therefore, the model that removes both modules (UFGraphFR-nB) is to implement both of the above implementations simultaneously. However, compared with the baseline model GPFedRec[1], it still has an additional UserMLP layer for further learning user preferences.
5.4.1 no Transformer
@@ 5.5 PLM & Modeling Effect(Q3)
  • LaBSE[46] The language-agnostic BERT sentence embedding encodes text into high-dimensional vectors. TheăEmbedding dimension is 768.
  • GTR-T5-Base,T5[47] This is a sentence transformers model: It maps sentences & paragraphs to a 768 dimensional dense vector space.
  • all-MiniLM-L6-v2,MiniLM This is a sentence transformers model: It maps sentences & paragraphs to a 384 dimensional dense vector space and can be used for tasks like clustering or semantic search.
@@ 6 Conclusion
Appendix A Previous submission to TNNLS for peer review
Recommended Decision:
  • Associate Editor: #1: Reject The overall evaluation of the current version is not so positive.
  • Associate Editor: #2: Reject The paper proposes a graph-based method for federated learning. The reviewers have found strong concerns about the lack of novelty, incomplete evaluations, incomplete references, and unsatisfactory presentations. The authors are suggested to substantially improve the work based on the reviewers’ detailed comments.
Individual Reviews:
  • Reviewer #1: This paper proposes UFGraphFR, a novel graph-enhanced federated recommendation framework that leverages user text descriptions to build user relationship graphs while ensuring privacy preservation. By introducing a joint embedding layer and Transformer-based sequence modeling, UFGraphFR captures user preferences without relying on explicit user interaction history or unique IDs. Additionally, a lightweight variant, UFGraphFR-Lite, reduces computational overhead by periodically updating the user graph, striking a balance between efficiency and performance. Experimental results on multiple benchmark datasets demonstrate the superior accuracy of UFGraphFR over state-of-the-art federated and centralized models. The proposed methods effectively enhance recommendation quality and privacy protection in federated learning scenarios.
    Pros:
    • 1
      The paper clearly outlines the limitations of existing methods and presents a well-defined solution.
    • 2
      The innovation level of this paper can be considered moderate to high. It introduces user text features and effectively integrates existing technologies, such as Transformers and GCNs.
    • 3
      The structure of this paper is relatively clear and well-organized.
    Cons:
    • 1
      The figure in the text does not seem so clear. For example, some of the elements in Figure 1. are vague.
    • 2
      There are some references that are not properly cited. Line 20 of Page 6, the references for FedMF are not shown correctly.
    • 3
      The authors should check the typos throughout this paper. Line 56 of Page 3, the word "in" is missing the "i".
    • 4
      Although the authors conducted ablation experiments for key components, maybe additional parameter ablation experiments is needed, such as $\lambda$ in the total loss function.
    • 5
      UFGraphFR-Lite periodically updates user diagrams, but the specific implementation details for the update frequency or the exact method of periodic updates do not appear to be provided in the paper.
    • 6
      The authors are encouraged to provide a detailed complexity analysis and discuss the scalability of the proposed method to large-scale datasets.
  • Reviewer #2: After going through the paper, although the references are cited accordingly which is pretty impressive the article requires a re-examination in terms of detailed explanation on novelty, the motivation behind in undertaking this research in general along with future scope improvements. This article looks a potential candidate for publication and will have be to be rejected if the above criteria are not met.
  • Reviewer #3: The paper introduces a UFGraphFR framework of Graph Federation Recommendation System based on User Text description Features, which focuses on using the text-attributed graph to enhance the recommendation while protecting the privacy. It tackles two main challenges: (1) How to build the graph without privacy exposure. (2) How to train the federated model based on the constructed text graph jointly.
    Strength:
    • 1
      The paper addresses a practical problem of federated recommendation with privacy protections, which is a significant topic in the community of federated learning.
    • 2
      Abundant Experimental results demonstrate the effectiveness of the proposed method.
    • 3
      The paper is readable and make it easy to follow.
    Weakness:
    • 1
      About novelty: Beyond the framework, there seem to be absence of original techniques and methods, with the approach primarily consisting of an amalgamation of existing techniques.
    • 2
      About clarity: the paper is filled with many typo errors. For example, the line 56 of page 3 and the line 20 for FedMF of page 6.
    • 3
      About experiments: (1) For datasets: It is unfair to compare only on small datasets and more large-scale datasets should be compared. (2) For evaluation: how to split the datasets? How many clients are participating during the training? Whether the data of each client is fairly divided? The author needs to provide a detailed explanation.
Appendix B Previous submission to EAAI for peer review
Final Decision: Reject
  • Reviewer #1:
    The manuscript presents a timely and well-motivated study on privacy-preserving recommendation systems through a federated learning framework. The proposed UFGraphFR model effectively combines semantic-aware graph construction with privacy safeguards, particularly through its use of joint-embedding strategies and a Transformer-based sequence encoder with differential privacy. The integration of on-device processing and graph neural networks is technically sound and addresses key limitations in prior work. However, a few minor points should be considered for improvement. The authors are encouraged to clarify the specific architecture and training details of the Joint-Embedding Layer to enhance reproducibility. Additionally, while the Transformer-based encoder is a strong component, its computational complexity on edge devices could be briefly discussed. A schematic summarizing the overall workflow would aid reader comprehension. Lastly, a short discussion on the scalability of UFGraphFR under varying user graph densities would further strengthen the manuscript. Overall, the paper is a valuable contribution and merits publication after minor revision.
  • Reviewer #2: The research paper titled "UFGraphFR: Graph Federation Recommendation System based on User Text description features" discusses a research area of significant interest to the scientific community. However, the paper is currently in a suboptimal state and requires a complete overhaul before a final decision can be made. Below are few suggestions to improve the quality of the manuscript:
    • 1
      It is suggested to briefly discuss the structure of subsequent sections in a single paragraph at the end of introduction section.
    • 2
      It is suggested to include a table of notations used in the subsequent analysis.
    • 3
      It is suggested to include a table in Section 2 discussing the pros and cons of popular techniques.
    • 4
      It is suggested to have the manuscript revised by an expert whose first language is English. Additionally, revise "Where" to lowercase "where" for each equation.
    • 5
      It is suggested to include a step-by-step algorithm outlining the proposed model.
    • 6
      It is suggested to perform statistical test such as Friedman test, Nemenyi test and Wilcoxon test.
    • 7
      It is suggested to include the DOI information for each research paper in the reference section. In addition, include at least three papers from 2025.
  • Reviewer #4: Rejected Not an interesting paper. just the abstract, literature, and methodology sections.